OpenStudy (anonymous):

Suppose y varies inversely with x. Write an equation for the inverse variation.

OpenStudy (anonymous):

27. y=7 when x=3 29. y= -9 when x=2

OpenStudy (kc_kennylau):

If y varies inversely with x, xy should be a constant.

OpenStudy (kc_kennylau):

Therefore you can set xy=k and find k

OpenStudy (mathmale):

Another way to look at this case of "inverse proportion" would be to write the general formula y=k/x. substituting the known values of x and y into this equation or Kenny's will give y ou the value of the constant of proportionality, k. Finally, write y=k/x, substituting the value of k that y ou've just found.

OpenStudy (anonymous):

"Inverse variation" basically means that one variable is varying in a way that it's in the form: \[y = \frac{k}{x}\] So use that to find out k and write the equation you've found out, which should be your answer. e.g, one answer for such a question will look like this: \[y = \frac{21}{x}\]